# LCSim: A Large-Scale Controllable Traffic Simulator

### v1

@@ C.2 Vehicle Behavior Analysis
We conducted statistical analysis on the dataset, focusing on metrics such as max acceleration, usual brake acceleration, velocity, relative distance, relative velocity, and time headway during the car following process, Figure 10 shows the results. This analysis allowed us to derive the driving behavior characteristics of vehicles in the Yizhuang area.
Figure 10: The analysis of SenseTime driving dataset.
Appendix D Multi-Style Reinforcement Learning
We constructed single-agent reinforcement learning experiments based on the Waymo traffic scenarios with our guided diffusive simulation to see the influence of styles of scenarios on policy learning.
D.1 Reinforcement Learning Setup
We constructed a reinforcement learning environment based on the validation set of the Waymo dataset. 4,400 scenarios are selected from the validation set and further divided into a training set containing 4,000 scenarios and a test set containing 400 scenarios. We trained a PPO [31] agent on the training set and evaluated its performance on the test set.
Observation Spec. Observation of the agent consists of two parts:
@@ D.1 Reinforcement Learning Setup
  • Scene Embedding: Embedding computed by scene encoder of the diffusion model with size of $[N_{h}]$, by applying cross attention to map polygons and agent states, this feature contains information about surrounding vehicles, road elements, and the vehicle’s own historical states. In this experiment, we use $N_{h}=128$ following the setup of the diffusion model.
@@
  • Route: We sampled the vehicle’s trajectory points within the next 1 second at a frequency of 10Hz and projected them into a relative coordinate system based on the vehicle’s current position and orientation. The shape of the route data is $[10,2]$, representing the reference path of the vehicle’s forward movement. If the vehicle behavior in the driving environment is generated by a diffusion model, then this path will be accumulated from the behavior sequences generated by the model for the vehicle.
Action Spec. We let the agent directly control the throttle and steering angle of the vehicle. The agent’s output is a two-dimensional vector with a range $[-1,1]$. This vector is multiplied by the maximum range of acceleration and steering angle, resulting in the final vehicle action. In this experiment, the maximum acceleration and steering angle of the vehicle are set to 6.0 and 0.3, respectively.
Rollout Setting. To let the agent explore every scenario in the training set, we randomly divided the 4000 scenes in the training set into 20 parts, each containing 200 different scenarios. We used 20 parallel threads to rollout episodes, with each thread pre-loading and pre-calculating map embeddings for 200 different training scenarios. During the rollout process, after the current episode ends, the environment automatically switches to the next scenario, and this cycle continues iteratively.
@@
Reward Function. Our goal is to make the vehicle progress along the given route while avoiding collisions and staying within the road. Therefore, we provide the following formula for the reward:
$R=R_{forward}+P_{collision}+P_{road}+P_{smooth}+R_{destination}.$ (4)
@@
  • $R_{destination}$: When an episode ends, we check if the vehicle has reached the destination of the given route, which means the distance to the endpoint of the route is within $2.5$ meters. If yes, the reward value is $10$; otherwise, it’s $-5$.
D.2 Multi-Style Environments Building
We build four kinds of environments with different driving styles using cost functions in Table 7:
  • The original Waymo driving environment, in this environment, vehicles base their actions on real trajectories from the Waymo driving logs.
  • The Waymo-style environment with diffusive simulation. This environment utilizes the diffusion model without guide functions, the vehicle behaviors are consistent with the Waymo dataset. With the diffusion model’s nature, it generates diverse vehicle trajectories under the same initial conditions, exposing the agent to a broader range of traffic scenarios during training.
  • The SenseTime-style environment with guided diffusive simulation. This environment follows the driving style observed in the SenseTime driving dataset, emphasizing a more "gentle" driving behavior compared to the Waymo-based environment. In this environment, we use cost functions on max acceleration with $acc_{max}=3m/s^{2}$, and on time headway with $thw_{target}=2.5s$.
  • The adversarial environment. This environment is implemented by guiding nearby vehicles closer to the vehicle controlled by the RL agent. For vehicles in front of or alongside the main vehicle, we guide their action generation with $dis_{target}=0$ to the main vehicle, thereby encouraging more sudden braking and cutting-in behaviors, increasing the aggressiveness of the environment.

### v2

@@ IV-B1 Private Driving Dataset
Our private driving dataset comprises about 400 hours of vehicle driving logs collected from vehicles in the Beijing Yizhuang area. The data is presented in a format similar to vehicle trajectories in WOMD. We conducted statistical analysis on the dataset, focusing on metrics such as acceleration, relative distance, and time headway during the car following process. This analysis allowed us to derive the driving behavior characteristics of vehicles in the Yizhuang area.
In Figure 7, we compare the behavioral characteristics of vehicles in our private dataset with those in WOMD. The comparison includes metrics such as vehicle acceleration, relative distance, and time headway during car following. Since a vehicle’s driving speed is often related to the specific driving environment (e.g., road congestion, lane speed limits) rather than the behavioral characteristics, we do not include speed in the comparison. It can be observed that there are significant differences between the behaviors of vehicles in these two datasets. Vehicles in the Yizhuang area exhibit a more "gentle" driving style, showing a preference for using smaller accelerations during start and brake. Additionally, they maintain larger relative distances and headway times during the following process compared to vehicles in the Waymo dataset.
As our model is trained on WOMD, without imposing any guidance on the generation process, the vehicle behavior characteristics in the diffusion-based simulation remain consistent with the Waymo dataset, as shown in Figure 6. By applying guide functions including max acceleration, relative distance, and time headway during the generation process, we can align the vehicle behavior characteristics produced by the diffusion-based simulation with those collected in our private driving dataset. The comparison of the two distributions can be seen in Figure 8. This demonstrates that our simulator can model vehicles with diverse driving styles, thereby providing traffic simulation environments with different driving styles.
Figure 7: The differences in vehicle behavior between WOMD and private datasets.
Figure 8: The comparison of vehicle behaviors between the private dataset and guided diffusion-based simulation.
IV-C Multi-Style RL Training
We construct a single-agent reinforcement learning environment based on WOMD with our guided diffusion-based simulation to further investigate the impact of traffic environments with different driving styles on driving policy learning. Below are the training settings and results:
Figure 9: Workflow of the RL agent.
IV-C1 Environment Settings
To validate the model’s capability in unseen scenarios, we construct a reinforcement learning environment based on the validation set of WOMD. We select 4,400 scenarios from the validation set and further divide them into a training set containing 4,000 scenarios and a test set containing 400 scenarios. We train a PPO [49] agent on the training set and evaluate its performance on the test set. As shown in Figure 9, we let the PPO agent control the self-driving car (SDC) marked in WOMD, the agent’s observation space consists of scene embedding computed by the scene encoder and a reference route for the vehicle. In different training environments, the route is either given by driving logs from the dataset or computed by motion plans generated by the diffusion model. The background vehicles are controlled by policies within our simulator. On the test set, we test four metrics of the agent: collision rate, off-road rate, average route progress rate, and scenario success rate. We also provide the average reward value per episode. The background vehicles of the test set act based on WOMD driving logs.
We create four distinct driving environments on the training set: In the first one, vehicles base their actions on real trajectories from WOMD driving logs. The second one utilizes the diffusion model without guide functions, which maintains consistency with WOMD in terms of vehicle behavior styles. With the diffusion model’s nature, it generates diverse vehicle motions under the same initial conditions, exposing the agent to a broader range of traffic scenarios during training. The third one follows the driving style observed in our private driving dataset, emphasizing a more "gentle" driving behavior compared to the WOMD-based environment. Furthermore, an adversarial driving environment is implemented by guiding nearby vehicles closer to the agent, creating a training scenario with a higher potential for collisions.
IV-C2 RL Settings
Observation of the agent consists of two parts:
@@ IV-C2 RL Settings
  • Scene Embedding: Embedding computed by scene encoder of the diffusion model with size of $[N_{h}]$, by applying cross attention to map polygons and agent states, this feature contains information about surrounding vehicles, road elements, and the vehicle’s own historical states. In this experiment, we use the size of $N_{h}$ following the setup of the diffusion model.
@@
  • Route: We sampled the vehicle’s trajectory points within the next one second at a frequency of 10Hz and projected them into a relative coordinate system based on the vehicle’s current position and orientation, representing the reference path of the vehicle’s forward movement. If the vehicle behavior in the driving environment is generated by a diffusion model, then this path will be accumulated from the motion plans generated by the model for the vehicle.
@@
Our goal is to make the vehicle progress along the given route while avoiding collisions and staying within the road. Therefore, we provide the following formula for the reward:
$R=R_{forward}+P_{collision}+P_{road}+P_{smooath}+R_{dest},$ (4)
@@
  • $R_{dest}$: When an episode ends, we check if the vehicle has reached the destination of the given route, which means the distance to the endpoint of the route is within $2.5$ meters. If yes, the reward value is $10$; otherwise, it’s $-5$.
TABLE V: Evaluation results of RL agents.
WOMD 15.71 ($\pm$ 0.56) 4.24 ($\pm$ 0.20) 84.98 ($\pm$ 0.53) 51.31 ($\pm$ 0.56) 7.41 ($\pm$ 0.10)
Diff w/o guide 12.06 ($\pm$ 0.28) 3.01 ($\pm$ 0.12) 84.76 ($\pm$ 0.35) 52.75 ($\pm$ 0.50) 8.59 ($\pm$ 0.14)
Figure 10: City-scale traffic scenario simulations in Jinan and Shenzhen.
IV-C3 Results
Table V presents the performance of agents trained in environments with different driving styles on the test set. Compared to agents trained on the original WOMD driving logs, those trained in diffusion-based simulation environments without guidance perform better across almost all metrics. This improvement is attributed to the diffusion-based simulations increasing the diversity of traffic scenarios in the training environment, enabling agents to learn more general and effective driving strategies. Conversely, agents trained in "gentler" environments perform poorly on the test set, as differences in background vehicle behavior between the training and test sets result in the agent’s driving strategies being ineffective in avoiding collisions. Additionally, the more passive driving style in the training environment’s reference routes leads to a lower route progress rate in the test set. Agents trained in adversarial environments excel at avoiding collisions, but their maneuvers to evade surrounding vehicles also result in a higher probability of driving off the road.
IV-D City-Scale Traffic Scenario Construction
We showcase the scalability of LCSim with city-scale simulations of real-world traffic scenarios in two metropolises.
IV-D1 Datasets
We use two vehicle trajectory datasets [54, 55, 56] each with one-day-long city-scale trajectories of the entire fleet recovered from daily urban traffic camera videos in Jinan and Shenzhen city. Both of the datasets involve over one million trajectories and one thousand square kilometers of urban area.
IV-D2 Settings
Compared with open-source driving datasets like WOMD, the trajectories recovered from traffic cameras are temporally sparser, where only the arrival time at road intersection is specified [54], which is thus taken as the travel schedule of each agent in LCSim with a series of trips between intersections with corresponding departure time and arrival time. We show that by simulating the vehicles given their schedules, using the arrival time at a specific position as goal point guidance, LCSim can effectively replicate real-world city-scale traffic scenarios.
IV-D3 Results
Figure 10 shows, in Jinan and Shenzhen, the spatial distribution of simulated road flow and speed, as well as the probability distribution of the arrival time error which is the deviation of the simulated arrival time of each trip compared with that of the ground truth trajectory data. As can be seen, the arrival time error is mainly distributed around zero with over 90% of trips having arrival time errors less than 20 seconds. LCSim also produces reasonable traffic conditions with coherence between the road network structure, flow, and speed.
IV-E Simulation Efficiency
TABLE VI: The attention mechanisms of scene encoder.
Simulator Metadrive Waymax CPU Waymax GPU LCSim w/o Diff LCSim w/ Diff
Time (s) 1.923 0.554 0.083 0.239 1.043
We compare LCSim and two baseline traffic simulators, MetaDrive and Waymax, we run scenarios in the validation set of WOMD with a length of 9s at 10Hz and compute the average simulation time per scenario. Table VI shows the result. We use an Intel(R) Xeon(R) Platinum 8462Y CPU and an NVIDIA GeForce RTX 4090 GPU for the simulation. And LCSim can achieve comparable results with other cpu-based simulators implemented in Python.
V Conclusion
We proposed LCSim, a large-scale, controllable diffusion-based traffic simulator. With an automated tool to construct traffic scenarios from multiple data sources, LCSim is capable of conducting large-scale traffic simulations. By integrating the diffusion-based motion planner and guide functions, LCSim can build traffic environments with diverse vehicle driving styles.
LCSim has two main limitations. Firstly, the simulator is implemented in Python using a single-threaded CPU, which limits its performance potential. Although parallel simulation using multiple processes is currently employed as a solution, it does not fundamentally address the issue. One potential approach to overcome this limitation is to develop a multi-threaded version of the simulator using C++ and deploy the Diffusion model in C++, which is a potential direction for future work. Secondly, the simulator currently provides visualization only from a top-down perspective and lacks the rendering of realistic perceptual data. Integrating the sensor data generation method into the simulation is one of the planned future developments to address this limitation.